# Sparse Autoencoders Reveal Universal Feature Spaces Across Large Language Models

### v3

@@ Abstract
We investigate feature universality in large language models (LLMs), a research field that aims to understand how different models similarly represent concepts in the latent spaces of their intermediate layers. Demonstrating feature universality allows discoveries about latent representations to generalize across several models. However, comparing features across LLMs is challenging due to polysemanticity, in which individual neurons often correspond to multiple features rather than distinct ones. This makes it difficult to disentangle and match features across different models. To address this issue, we employ a method known as dictionary learning by using sparse autoencoders (SAEs) to transform LLM activations into more interpretable spaces spanned by neurons corresponding to individual features. After matching feature neurons across models via activation correlation, we apply representational space similarity metrics on SAE feature spaces across different LLMs. Our experiments reveal significant similarities in SAE feature spaces across various LLMs, providing new evidence for feature universality  [^1] Code for experiments is available at: https://github.com/wlg1/univ_feat_geom .
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable capabilities across a wide range of natural language tasks (Bubeck et al., 2023; Naveed et al., 2024; Liu et al., 2023). However, understanding how these models represent and process information internally remains a significant challenge (Bereska & Gavves, 2024). A key question in this domain is whether different LLMs learn similar internal representations, or if each model develops its own unique way of encoding linguistic knowledge. This question of feature universality is crucial for several reasons: it impacts the generalizability of interpretability findings across models, it could accelerate the development of more efficient training techniques, and it may provide insights into safer and more controllable AI systems (Chughtai et al., 2023b; Gurnee et al., 2024; Sharkey et al., 2024; Bricken et al., 2023).
Comparing features across LLMs is inherently difficult due to the phenomenon of polysemanticity, where individual neurons often correspond to multiple features rather than distinct ones (Elhage et al., 2022). This superposition of features makes it challenging to disentangle and match feature representations across different models. To address these challenges, we propose a novel approach leveraging sparse autoencoders (SAEs) to transform LLM activations into more interpretable spaces before matching features. SAEs allow us to decompose the complex, superposed representations within LLMs into features that correspond to distinct semantic concepts that are easier to analyze and compare across models (Makhzani & Frey, 2013; Cunningham et al., 2023; Bricken et al., 2023). We then apply representational space similarity metrics to these SAE features, enabling a rigorous quantitative analysis of feature universality across different LLMs (Klabunde et al., 2023). Our methodology involves several key steps: obtaining SAEs trained on the activations of multiple LLMs to extract interpretable features, developing and applying representational similarity metrics tailored for high-dimensional SAE feature spaces, and conducting extensive experiments comparing SAE features across models with varying architectures, sizes, and training regimes.
To verify the effectiveness of our approach, we present a comprehensive set of experiments and results. Our experimental approach involves solving two issues for representational space similarity measurement: the permutation issue, which involves aligning individual neurons, and the rotational issue, which involves comparing two spaces which have different basis that are rotationally-invariant. Additionally, we present an approach to filter out feature matches with low semantically-meaningful similarity.
We demonstrate high correlations between individual SAE features extracted from distinct models, providing evidence for feature universality. Furthermore, we show that semantically meaningful subspaces (e.g., features related to specific concepts, such as Emotions) exhibit even higher similarity across models. We also analyze how feature similarity varies across different layers of LLMs, revealing patterns in how representations evolve through the model depth. Lastly, we examine our results to note the differences which are not captured by SAEs. While we discover evidence that LLMs learn weakly universal features, and that SAEs can capture both these features and these relations, we also find that different SAEs do not consistently learn the same universal features captured by LLMs, a finding that is supported by previous work (Leask et al., 2025). Thus, our research guides future work to improve upon methods that can better capture universal representations of features (Thasarathan et al., 2025), and contributes to ongoing research that aims to improve SAEs (Engels et al., 2024b).
This work opens up several exciting avenues for future research. Exploring whether universal features can be leveraged to improve transfer learning or model compression techniques may lead to more efficient training paradigms (Yosinski et al., 2014; Hinton et al., 2015). Additionally, examining the implications of feature universality for AI safety and control, particularly in identifying and mitigating potentially harmful representations, could contribute to the development of safer AI systems (Hendrycks et al., 2023). By shedding light on the internal representations of LLMs, our research contributes to the broader goal of developing more interpretable, efficient, and reliable AI systems (Barez et al., 2023). Understanding the commonalities and differences in how various models encode information is a crucial step toward building a more comprehensive theory of how large language models process and generate human language. Our key contributions include:
@@
    A novel framework for comparing LLM representations using SAEs and similarity metrics.
@@
    Empirical evidence supporting the existence of universal features across diverse LLMs.
@@
    A quantitative analysis of how feature similarity varies across layers and semantic domains.
  4. 4.
    A step-by-step approach to compare feature weight representations by paired features that differs from previous methods which compare feature activations.
@@
Figure 1: We train SAEs on LLMs, and then compare their SAE feature space similarity using metrics such as Singular Value Canonical Correlation Analysis (SVCCA).
2 Background
@@
Feature Universality. Previous studies have explored the concept of universality using various definitions Bricken et al. (2023); Chughtai et al. (2023a). In this paper, we examine universality through the lens of two definitions: analogous features (Olah et al., 2020), and representational universality (Gurnee et al., 2024). We define correlated features as features which are correlated in terms of activation similarity, and thus activate highly on the same tokens. We measure the representational universality of feature spaces by comparing their correlated feature pairs. Correlated features that reside in spaces with notable representational similarity are deemed as analogous. Further discussion of these definitions is provided in Appendix §H.
@@
Sparse Autoencoders. Sparse autoencoders (SAEs) are a type of autoencoder designed to learn efficient, sparse representations of input data by imposing a sparsity constraint on the hidden layer (Makhzani & Frey, 2013). The network takes in an input $\mathbf{x}\in\mathbb{R}^{n}$ and reconstructs it as an output $\mathbf{\hat{x}}$ using the equation $\mathbf{\hat{x}}=\mathbf{W}^{\prime}\sigma(\mathbf{W}\mathbf{x}+\mathbf{b})$, where $\mathbf{W}\in\mathbb{R}^{h\times n}$ is the encoder weight matrix, $\mathbf{b}$ is a bias term, $\sigma$ is a nonlinear activation function, and $\mathbf{W}^{\prime}$ is the decoder matrix, which often uses the transpose of the encoder weights. SAE training aims to both encourage sparsity in the activations $\mathbf{h}=\sigma(\mathbf{W}\mathbf{x}+\mathbf{b})$ and to minimize the reconstruction loss $\left\|\mathbf{x}-\hat{\mathbf{x}}\right\|_{2}^{2}$.
The sparsity constraint encourages a large number of the neuron activations to remain inactive (close to zero) for any given input, while a few neurons, called feature neurons, are highly active. Recent work has replaced the regularization term by using the TopK activation function as the nonlinear activation function, which selects only the top K feature neurons with the highest activations (Gao et al., 2024), zeroing out the other activations. The active feature neurons tend to activate only for specific concepts in the data, promoting monosemanticity. Therefore, since there is a mapping from LLM neurons to SAE feature neurons that "translates" LLM activations to features, this method is a type of dictionary learning (Olshausen & Field, 1997).
@@
SAE Feature Weight Spaces. By analyzing UMAPs of feature weights from an SAE trained on a middle residual stream layer of Claude 3 Sonnet, researchers discovered SAE feature spaces organized in neighborhoods of semantic concepts, allowing them to identify subspaces corresponding to concepts such as biology and conflict (Templeton et al., 2024). Additionally, these concepts appear to be organized in hierarchical clusters, such as "disease" clusters that contain sub-clusters of specific diseases like flus. In Section §4.3, we measure the extent of similarity of these semantic feature spaces across LLMs.
@@
Representational Space Similarity. Representational similarity measures typically compare neural network activations by assessing the similarity between activations from a consistent set of inputs (Klabunde et al., 2023). In this paper, we take a different approach by comparing the representations using the decoder weight matrices $W^{\prime}$ of SAEs, where the columns correspond to the feature neurons. We calculate a similarity score $m(W,W^{\prime})$ for pairs of representations $W$ in SAE $A$ and $W^{\prime}$ from SAE $B$. We obtain two scores via the following two representational similarity metrics:
@@
Singular Value Canonical Correlation Analysis (SVCCA): Canonical Correlation Analysis (CCA) seeks to identify pairs of canonical variables, $\mathbf{u}_{i}$ and $\mathbf{v}_{i}$, from two sets of variables, $\mathbf{X}\in\mathbb{R}^{n\times d_{1}}$ and $\mathbf{Y}\in\mathbb{R}^{n\times d_{2}}$, that are maximally correlated with each other (Hotelling, 1936). Singular Value Canonical Correlation Analysis (SVCCA) (Raghu et al., 2017) enhances CCA by first applying Singular Value Decomposition (SVD) to both $\mathbf{X}$ and $\mathbf{Y}$, which can be expressed as:
@@
where $\mathbf{U}_{X}$ and $\mathbf{U}_{Y}$ are the matrices containing the left singular vectors (informative directions), and $\mathbf{S}_{X}$ and $\mathbf{S}_{Y}$ are diagonal matrices containing the singular values. By projecting the original data onto these informative directions, noise is reduced. SVCCA then applies CCA on the transformed datasets, $\mathbf{U}_{X}$ and $\mathbf{U}_{Y}$, resulting in correlation scores between the most informative components. These scores are averaged to obtain a similarity score.
Representational Similarity Analysis (RSA): Representational Similarity Analysis (RSA) (Kriegeskorte et al., 2008) works by first computing, for each space, a Representational Dissimilarity Matrix (RDM) $\mathbf{D}\in\mathbb{R}^{n\times n}$, where each element of the matrix represents the dissimilarity (or similarity) between every pair of data points within a space. The RDM essentially summarizes the pairwise distances between all possible data pairs in the feature space. A common distance metric used is the Euclidean distance. After an RDM is computed for each space, a correlation measurement like Spearman’s rank correlation coefficient is applied to the pair of RDMs to obtain a similarity score. RSA has been applied to measure relational similarity of stimuli across brain regions and of activations across neural network layers (Klabunde et al., 2023).
SVCCA vs RSA. SVCCA focuses on the comparison of entire vector spaces by aligning the principal components (singular vectors) of two representation spaces and then computing the canonical correlations between them. In other words, it identifies directions in the space that carry the most variance and checks how well these align between two models or layers. On the other hand, RSA is primarily concerned with the structure of relationships, or relational structure, between different data points (like stimuli) within a representation space. It allows us to measure how well relations such as king-queen transfer across models. Thus, SVCCA focuses on comparing the aligned subspaces of two representation spaces, while RSA is more suitable for analyzing the structural similarity between representations at the level of pairwise relationships.
@@ 3 Methodology
Representational similarity is important for capturing the across-model generalizations of both: 1) Feature Spaces, which are spaces spanned by groups of features, and 2) Feature Relations. We compare SAEs trained on layer $A_{i}$ from LLM $A$ and layer $B_{j}$ from LLM $B$. This is done for every layer pair. To compare latent spaces using our selected metrics, we have to solve both permutation and rotational alignment issues. For the permutation issue, we have to find a way to pair neuron weights. Typically, these metrics are performed by pairing activation vectors from two models according to common input instances. For example, input $I_{k}$ is passed through models $A$ and $B$, and then a vector of $N_{1}$ neuron activations from model $A$ is compared to a paired vector of $N_{2}$ activations from model $B$. The common input instance creates individual mappings from one space to another, which is required for metrics like SVCCA and RSA (Klabunde et al., 2023).
In our paper, we do not pair activation vectors by input instances, but we pair the feature neuron weights themselves. This means we have to find individual pairings of "similar" neurons in SAE $A$ to SAE $B$. However, we do not know which features map to which features; the order of these features are permuted in the weight matrices, and some features may not have a corresponding partner. Therefore, we have to pairwise match them via a correlation metric. For the rotational issue, models typically develop their own basis for latent space representations, so features may be similar relation-wise across spaces, but not rotation-wise due to differing basis. To solve this issue, we employ rotation-invariant metrics.
Assessing Scores with Baseline. We follow the approach of Kriegeskorte et al. (2008) to randomly pair features to obtain a baseline score. We compare the score of the features paired by correlation (which we refer to as the "paired features") with the average score of $N$ runs of randomly paired features to obtain a p-value score. If the p-value is less than 0.05, the metric suggests that the similarity is statistically meaningful.
In summary, the steps to carry out our similarity comparison experiments are given below, and steps 1 to 3 are illustrated in Figure 19:
@@
    For the permutation issue: Get activation correlations for every feature pair from two decoder SAEs for similar LLMs. Pair each feature from model $A$ with its highest activation correlated feature from model $B$.
@@
    For the rotational issue: Apply various rotation-invariant metrics to get a Paired Score.
@@
    Using the same metrics, obtain the similarity scores of $N$ random pairings to estimate a null distribution. Obtain a p-value of where the paired score falls in the null distribution to determine statistical significance.
Semantic Subspace Matching Experiments. For these experiments, we first identify semantically similar subspaces, and then compare their similarity. This is summarized by the following steps:
  1. 1.
    We find a group of related features in each model by searching for features which highly activate on the top 5 samples’ tokens that belong to a concept category. For example, the concept of "emotions" contains keywords such as "rage". Since the correlations are calculated by tokens, we use single-token keywords. If a keyword from a concept appears at least once, that feature is included. The keywords in each category are given in Appendix L.
  2. 2.
    We obtain mappings between these subsets of features using max activation correlation.
  3. 3.
    We characterize the relationships of features within these groups (eg. pairwise differences between features), and compare these relationships across models using similarity metrics.
Assessing Scores with Baseline. We use two types of tests which compare the paired score to a null distribution. Each test examines that the score is rare under a certain null distribution assumption:
@@
    Test 2: We compare the score of a semantic subspace mapping to a mean score of mappings between randomly selected feature subsets of the same sizes as the semantic subspaces. This uses a correlation matrix only between the two subsets, not between the entire space. This shows that the high similarity score does not hold for any two subspaces of the same size.
Metrics for the Permutation Issue: These metrics measure individual, local similarity.
Highest Activation Correlation. We take the activation correlations between every feature, following the approach of Bricken et al. (2023). We pass a dataset of samples through both models. Then, for each feature in model $A$, we find its activation correlation with every feature in model $B$, using tokens as the common instance pairing. Lastly, for each feature in model $A$, we find its highest correlated feature in model $B$ (as this allows for many-to-1 mappings, this is termed 1A-to-ManyB; the other way around is ManyA-to-1B). We pair these features together when conducting our experiments. We refer to scores obtained by pairing via highest activation correlation as Paired Scores.
Filter by Top Tokens. We notice that there are many "non-concept mappings" with "non-concept features" that have top 5 token activations on end-of-text / padding tokens, spaces, new lines, and punctuation; these non-concept mappings greatly bring our scores down as they are not accurately pairing features representing concepts, as we describe further in Appendix K. By filtering max activation correlation mappings to remove non-concept mapping features, we significantly raise the similarity scores. We also only keep mappings that share at least one keyword in their top 5 token activations. The list of non-concept keywords is given in the Appendix K.
Filter Many-to-1 Mappings. We find that some of these mappings are many-to-1, which means that more than one feature maps to the same feature. We found that removing many-to-1 mappings slightly increased the scores, and we discuss possible reasons why this may occur in Appendix J. However, the scores still show high similarity even with the inclusion of many-to-1 mappings, as seen in Figures 2. We show only scores of 1-to-1 mappings (filtering out the many-to-1 mappings) in the main text, and show scores of many-to-1 mappings in Appendix J. We use features from the SAEs of Pythia-70m and Gemma-1-2B as "many" features which map onto "one" feature in the SAEs of Pythia-160m and Gemma-2-2B.
Metrics for the Rotational Issue: Rotation-invariant metrics measure global similarity after feature pairing. We apply metrics which measure: 1) How well subspaces align using SVCCA, and 2) How similar feature relations like king-queen are using RSA. We set the inner similarity function using Pearson correlation the outer similarity function using Spearman correlation, and compute the RDM using Euclidean Distance.

### v4

# Quantifying Feature Space Universality Across Large Language Models via Sparse Autoencoders

@@ Abstract
The Universality Hypothesis in large language models (LLMs) claims that different models converge towards similar concept representations in their latent spaces. Providing evidence for this hypothesis would enable researchers to exploit universal properties, facilitating the generalization of mechanistic interpretability techniques across models. Previous works studied if LLMs learned the same features, which are internal representations that activate on specific concepts. Since comparing features across LLMs is challenging due to polysemanticity, in which LLM neurons often correspond to multiple unrelated features, rather to than distinct concepts, sparse autoencoders (SAEs) have been employed to disentangle LLM neurons into SAE features corresponding to distinct concepts. In this paper, we introduce a new variation of the universality hypothesis called Analogous Feature Universality: we hypothesize that even if SAEs across different models learn different feature representations, the spaces spanned by SAE features are similar, such that one SAE space is similar to another SAE space under rotation-invariant transformations. Evidence for this hypothesis would imply that interpretability techniques related to latent spaces, such as steering vectors, may be transferred across models via certain transformations. To investigate this hypothesis, we first pair SAE features across different models via activation correlation, and then measure spatial relation similarities between paired features via representational similarity measures, which transform spaces into representations that reveal hidden relational similarities. Our experiments demonstrate high similarities for SAE feature spaces across various LLMs, providing evidence for feature space universality.  [^1] Code for experiments is available at: https://github.com/wlg1/univ_feat_geom
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable capabilities across a wide range of natural language tasks (Bubeck et al., 2023; Naveed et al., 2024). However, understanding how these models represent and process information internally remains a significant challenge (Bereska and Gavves, 2024). The Universality Hypothesis claims that different models converge towards similar concept representations in their latent spaces (Huh et al., 2024). Providing evidence for this hypothesis would enable researchers to exploit universal properties related to features, providing new ways for generalizating mechanistic interpretability techniques across models (Sharkey et al., 2024), and accelerating research towards safer and more controllable AI systems (Hendrycks et al., 2023).
If LLMs learn similar internal concept representations, called features, it is vital is to understand which measures they are similar under, and to quantify the extent of these similarities (Chughtai et al., 2023; Gurnee et al., 2024). Since comparing features across LLMs is challenging due to polysemanticity, in which individual neurons often correspond to multiple unrelated features, rather than to distinct concepts, sparse autoencoders (SAEs) have been employed to disentangle polysemantic LLM neurons into SAE features corresponding to distinct semantic concepts that are easier to analyze and compare across models (Cunningham et al., 2023). Previous works have found that while SAEs learn similar features, they also learn a notable percentage of different features that are idiosyncratic to specific SAEs (Leask et al., 2025; Paulo and Belrose, 2025). Even so, SAEs have been found to learn correlated features which activate on the same concepts across models (Bricken et al., 2023). Moreover, semantically similar SAE feature arrangements are found across SAEs trained on different models, such as "months" features arranged circularly in GPT-2, Mistral-7B, and Llama-3 8B (Engels et al., 2025; Leask et al., 2024). Research has also found other types of feature arrangements (Park et al., 2024a; Li et al., 2025; Ye et al., 2024). These findings suggest investigating a new way to think of universality from the perspective of feature relations.
In this paper, we introduce a new variation of the universality hypothesis called Analogous Feature Universality: we hypothesize that even if SAEs across different models learn different feature representations, the spaces spanned by SAE features are similar, such that one SAE space is similar to another SAE space under certain rotation-invariant transformations. Evidence supporting this hypothesis would imply that latent spaces may have similar geometric arrangements of semantic subspaces (Templeton et al., 2024), suggesting that features, and techniques related to feature spaces (e.g. steering vectors), may be transferred across models via particular transformations. Previous works transfer features across models (Merullo et al., 2023), but there is a lack of research explaining why this can be done.
To investigate this hypothesis, we develop a novel approach that involves two main steps: (1) first, we pair SAE features across different models via activation correlation, and (2) afterwards, we measure the similarity of spatial relations between paired features using representational space similarity measures, which transform spaces into representations that reveal hidden relational similarities (Raghu et al., 2017; Kriegeskorte et al., 2008). This approach addresses a different research question compared to previous representational similarity papers (Kornblith et al., 2019; Klabunde et al., 2024, 2023; Sucholutsky et al., 2023), which did not measure the relationships among features, but among input data samples that were already known to correspond to the “same point” in input data space. In contrast, to investigate our hypothesis, our approach differs from previous approaches as it introduces step 1, which allows us to assess if we can use features, not input samples, to determine if representations are the same across models. Thus, we are applying representational similarity based on features in weight space, instead of based on input data samples in activation space.
To better describe why this approach works for testing our hypothesis, we can consider the following analogy: given three points [1, 2, 3] on a right triangle, we can think of each point like a feature. One model may label these points as [A, B, C], with each position being a column of a matrix, and another model may label them as [X, Y, Z]. However, these points are not ordered the same in each matrix, so we may find them in the order of [B, C, A] and [Z, X, Y]. But if we rearrange them into pairs “close enough to” [1, 2, 3] by estimating which points they correspond to on a right triangle, then we can align the two representations, and measure if the relationship among them—a right triangle—is also similar. This necessitates step 1 in our approach: consistent high scores across models can only be found if these features are paired correctly.
Our experiments demonstrate high similarities in SAE feature spaces between middle layers across multiple LLMs, providing evidence for feature space universality. Furthermore, we show that semantically meaningful subspaces exhibit notable similarity across models.
Our key contributions include:
@@
    Empirical evidence for similar SAE feature space representations across diverse LLMs.
@@
    A novel approach to study similar feature spaces by comparing weight representations, instead of activation representations, via paired features.
@@
    A quantitative analysis of how feature similarity varies by semantic subspaces (such as "Emotions").
@@
Figure 1: We train SAEs on LLMs, and then compare their SAE feature space similarity using measures such as Singular Value Canonical Correlation Analysis (SVCCA).
2 Background and Definitions
@@
Feature Universality. We define a feature as an internal representation that activates on a concept (Park et al., 2024b). To quantify the extent of feature space universality, we utilize the following definitions that build on terms defined in previous works: We define correlated features as features which are correlated in terms of activation similarity, and thus activate highly on the same tokens (Bricken et al., 2023). We measure the representational universality of feature spaces by applying representational similarity measures on a set of correlated feature pairs (Gurnee et al., 2024). Correlated features that reside in spaces with statistically significant representational similarity (as measured by the main approach discussed in Section §3) are labeled as analogous features (Olah et al., 2020). These similarity measures also measure how much analogical similarity, relative to a set of pairings and transformations, each feature has to its paired counterpart. Further discussion of these definitions is provided in Appendix §H.
We do not measure how well models capture true features, which are atomic ground truth features that models may converge towards capturing (Chughtai et al., 2023). Our hypothesis is that even if SAEs do not capture the same exact features, their features can be matched as analogous features that lie in spaces which can be transformed from one to another.
Sparse Autoencoders. Superposition is a phenomenon in which a model, in response to the issue of having to represent more features than it has parameters, learns feature representations distributed across many parameters (Elhage et al., 2022). This causes its parameters, or neurons, to be polysemantic, which means that each neuron is involved in representing more than one feature, making it difficult to compare features across different LLMs. To address this issue, Sparse Autoencoders (SAEs) have been applied to disentangle an LLM’s polysemantic neuron activations into monosemantic "feature neurons", which are encouraged to represent an isolated concept.
@@
SAEs are a type of autoencoder that learn sparse representations of input data by imposing a sparsity constraint on the hidden layer (Makhzani and Frey, 2013; Cunningham et al., 2023). An SAE takes in LLM layer activations $\mathbf{x}\in\mathbb{R}^{n}$ and reconstructs $\mathbf{x}$ as output $\mathbf{\hat{x}}=\mathbf{W}^{\prime}\sigma(\mathbf{W}\mathbf{x}+\mathbf{b})$, where $\mathbf{W}\in\mathbb{R}^{h\times n}$ is the encoder weight matrix, $\mathbf{b}$ is a bias term, $\sigma$ is a nonlinear activation function, and $\mathbf{W}^{\prime}$ is the decoder matrix, which often uses the transpose of the encoder weights. SAE training aims to both encourage sparsity in the activations $\mathbf{h}=\sigma(\mathbf{W}\mathbf{x}+\mathbf{b})$ and to minimize the reconstruction loss $\left\|\mathbf{x}-\hat{\mathbf{x}}\right\|_{2}^{2}$. The sparsity constraint encourages a large number of SAE neuron activations to remain inactive for any given input, while a few neurons, called feature neurons, are highly active. The active feature neurons tend to activate only for specific concepts in the data, promoting monosemanticity. Therefore, since there is a mapping from LLM neurons to SAE feature neurons that "translates" LLM activations to features, this method is a type of dictionary learning (Olshausen and Field, 1997). Several works make architectural improvements to SAEs (Gao et al., 2025; Rajamanoharan et al., 2024a, b), and analyze its ability to capture LLM features (Chanin et al., 2024).
@@
SAE Feature Weight Spaces. By analyzing feature weights UMAPs of an SAE trained on a middle residual stream layer of Claude 3 Sonnet, researchers discovered feature spaces organized in neighborhoods of semantic concepts, and identified subspaces corresponding to concepts such as "biology" and "conflict" (Templeton et al., 2024). Additionally, these concepts appear to be organized in hierarchical clusters, such as "disease" clusters that contain sub-clusters of specific diseases like flus. In Section §4.3, we measure the extent of similarity of semantic feature spaces across LLMs.
@@
Representational Space Similarity. Representational similarity measures typically compare neural network activations by assessing the similarity between activations from a consistent set of inputs (Klabunde et al., 2023). In this paper, we take a different approach by comparing the representations using the SAE decoder weight matrices $W^{\prime}$, whose columns correspond to feature neurons. We calculate a similarity score $m(W_{X}^{\prime},W_{Y}^{\prime})$ for pairs of representations from SAEs $X$ We obtain two scores via the following representational similarity measures  [^2] Appendix I discusses details on how these measures are applied on decoder weight matrices.:
@@
Singular Value Canonical Correlation Analysis (SVCCA): Given two sets of variables, $\mathbf{X}\in\mathbb{R}^{n\times d_{1}}$ and $\mathbf{Y}\in\mathbb{R}^{n\times d_{2}}$, Canonical Correlation Analysis (CCA) seeks to identify pairs of linear transformations $\mathbf{a}$ and $\mathbf{b}$ that maximize the correlation between $\mathbf{X}a$ and $\mathbf{Y}b$ (Hotelling, 1936). Singular Value Canonical Correlation Analysis (SVCCA) (Raghu et al., 2017) enhances CCA by first applying Singular Value Decomposition (SVD) to both $\mathbf{X}$ and $\mathbf{Y}$ to obtain the left singular matrices $\mathbf{U}_{X}$ and $\mathbf{U}_{Y}$ which form orthonormal bases for the column space of their matrices:
@@
$\mathbf{S}_{X}$ and $\mathbf{S}_{Y}$ are diagonal matrices containing the singular values, and $\mathbf{V}_{X}^{T}$ and $\mathbf{V}_{Y}^{T}$ contain the right singular vectors. Dimensionality reduction is applied by truncating low variance directions, which reduces noise, to obtain $\mathbf{U}_{X,k}$ and $\mathbf{U}_{Y,k}$. These represent the principal directions of variance (most informative directions).
Then, CCA is applied on $\mathbf{U}_{X,k}$ and $\mathbf{U}_{Y,k}$ to find linear transformations $a$ and $b$ that maximize the Pearson correlation between $\mathbf{U}_{X,k}a$ and $\mathbf{U}_{Y,k}b$, such that $a$ and $b$ are the canonical directions in each space that are maximally correlated. The SVCCA score is the mean of the top canonical correlations. In other words, it measures how well linear relationships $\mathbf{U}_{X,k}a$ and $\mathbf{U}_{Y,k}b$ align.
Representational Similarity Analysis (RSA): Representational Similarity Analysis (RSA) (Kriegeskorte et al., 2008) has been applied to measure relational similarity of stimuli across brain regions and of activations across neural network layers (Klabunde et al., 2023). First, for each space $\mathbf{X}\in\mathbb{R}^{n\times d_{1}}$ and $\mathbf{Y}\in\mathbb{R}^{n\times d_{2}}$, RSA constructs a Representational Dissimilarity Matrix (RDM) $\mathbf{D}\in\mathbb{R}^{n\times n}$, where each element of the matrix represents the dissimilarity (or similarity) between every pair of data points within a space. The RDM essentially summarizes the pairwise distances between all possible data pairs in the feature space. For space $\mathbf{X}$, the RDM has entries:
$D_{ij}(X)=\|x_{i}-x_{j}\|_{2},$
where $x_{i}$ and $x_{j}$ are row vectors (e.g., features). A common distance metric used is the Euclidean distance. After an RDM is computed for each space, a correlation measurement like Spearman’s rank correlation coefficient is applied to the pair of RDMs to obtain a similarity score.
@@ 3 Methodology
We compare SAEs trained on layer $A_{i}$ from LLM $A$ and layer $B_{j}$ from LLM $B$. This is done for every layer pair. To compare latent spaces using our selected measures, we have to solve both permutation and rotational alignment issues. For the permutation issue, we have to find a way to pair neuron weights. This was not an issue in previous representational similarity studies because activations were already paired by input data samples. (Klabunde et al., 2023). For our case, we do not pair activation vectors by input instances, but pair by feature neuron weights. This means we have to find neuron pairings in SAE $S_{A}$ to SAE $S_{B}$ based on individual local similarity. However, we do not know which features map to which features; the orderings of these features are permuted in the weight matrices, and some features may not have a corresponding partner. Therefore, we pair them via a correlation metric; a set of pairings is a mapping. For the rotational issue, models learn their own distinct basis for latent space representations, so features may be similar relation-wise across spaces, but not rotation-wise due to differing bases. To solve this, we employ rotation-invariant measures.
Assessing Scores with Baseline. We follow the approach of Kriegeskorte et al. (2008) to randomly pair features to obtain a baseline score. We compare the score of the features paired by correlation (which we refer to as the "paired features") with the average score of $N$ runs of randomly paired features to obtain a p-value score. If the p-value is less than 0.05, the measure suggests that the similarity is statistically meaningful. In summary, the steps to carry out our similarity comparison experiments are given below, and steps 1 to 3 are illustrated in Figure 19:
@@
    For the permutation issue: Get activation correlations for feature pairs from SAE decoder weights. Pair features from SAE $A$ with its highest activation correlated feature from SAE $B$.
@@
    For the rotational issue: Apply rotation-invariant measures to get a Paired Score.
@@
    Using the same measures, obtain the similarity scores of $N$ random pairings to estimate a null distribution. Obtain a p-value of where the paired score falls in the null distribution to determine statistical significance.
Solving the Permutation Issue: Due to learned parameter variations specific to each LLM and SAE, we do not expect every feature in a SAE to have a corresponding feature that is "close enough" in another SAE. Else, we would forcibly pair features without corresponding partners, yielding low scores due to "bad pairings". Thus, we compare SAE feature subsets.
Highest Activation Correlation. We take the activation correlations between every feature, following the approach of Bricken et al. (2023). We pass a dataset of samples through both models. Then, for each feature in SAE $A$, we find its activation correlation with every feature in SAE $B$, using tokens as the common instance pairing. Lastly, for each feature in SAE $A$, we find its highest correlated feature in SAE $B$ (as this allows for many-to-1 mappings, this is termed 1A-to-ManyB; the other way around is ManyA-to-1B). We pair these features together when conducting our experiments. We refer to scores obtained by pairing via highest activation correlation as Paired Scores.
Filter by Top Tokens. We notice some pairings with "non-concept features" that have top 5 token activations on end-of-text / padding tokens, spaces, new lines, and punctuation; these non-concept pairings greatly reduce scores possibly due to being "bad pairings", as we describe further in Appendix K. By filtering out non-concept pairings, we significantly raise the similarity scores. We also only keep pairings that share at least one keyword in their top 5 token activations. The list of non-concept keywords is given in the Appendix K.
Filter Many-to-1 Mappings. We find that some mappings are many-to-1, meaning more than one feature maps to the same feature. We found that removing many-to-1 pairings slightly increased the scores, and we discuss possible reasons why this occurs in Appendix J. However, the scores still show high similarity even with the inclusion of many-to-1 pairings. We show scores of 1-to-1 pairings (filtering out many-to-1 pairings) in the main text, and show scores of many-to-1 pairings in Appendix J. We choose "many" features from the SAEs of the smaller LLM in a model pair, and find similar results if we choose from SAE of the larger LLM.
[…]
Semantic Subspace Matching Experiments. For these experiments, we first identify semantically similar subspaces, and then compare their similarity. We find a group of related features in each model by searching for features which highly activate on the top 5 samples’ tokens that belong to a concept category. For example, the "emotions" concept contains keywords like "rage".
Baseline Tests Types. We use two types of tests which compare the paired score to a null distribution. Each test examines that the score is rare under a certain null distribution assumption.
@@
    Test 2: We compare the score of a semantic subspace mapping to a mean score of mappings between randomly selected feature subsets of the same sizes as the semantic subspaces. This shows that the high similarity score does not hold for any two subspaces of the same size.